OpenStudy (anonymous):

Write a recursive definition for the sequence 8, 6, 4, 2, … (1 point) I Think Its a1 = 8; an= an–1 − 2 But I Just Wanna Make Sure!

OpenStudy (anonymous):

\[a_1=8\\ a_2=6=8-2=a_1-2\\ \vdots\\ a_n=a_{n-1}-2\] Yes, it works out.